# Literature Review, Methodology, Software, Data curation, Implementation, Experimental analysis, Writing - Original draft preparation

### v1

@@ 5 Explainability Analysis
LIME and Transformer Interpret both showed how the fine-tuned DistilBERT model arrives at its prediction on phishing email detection by approximating the model’s behaviour on every word of the email text. Lime and Transformer Interpret provided accurate results by answering the “Why” model predicting the selected text is a “Safe Email” or “Phishing Email”. Based on the explainability, those explainable techniques generate word attributions by providing positive and negative numbers on each word. Those numbers showed which word was more specifically reasonable for the predicted class. By using those numbers the explainable model visualizes the whole text by highlighting the words that scored positive and as well as negative on the prediction.
For this explainability analysis in LIME explainer, we initialized the number of features (words) 15 for the dimensionality reduction. If the number of words of a selected test is very large that time the computational time will increase as well as it increases the dimensionality which also can be a reason why the systems crashed. Limiting this number of features makes the explanation process computationally more feasible and provides a clearer and more concise picture of which words in the case of text are influential in the model’s prediction for a particular instance. Fig. 8 shows the model explainability using LIME and Table  7 and Table  8 show the positive and negative weights associated with words that help users understand phishing email detection.
(a) Explanation for the first text
(b) Explanation for the second text
Figure 8: LIME Explanation for two Phishing Email Texts
Table 7: The positive and negative weights using LIME for first text
Positive Numbers Negative Numbers
Word Value Word Value
‘low’ 1.760212017947662 ‘neonate’ -0.1267345497528774
‘price’ 1.5916451492409904 ‘3107’ -0.013956063314091143
‘software’ 1.2586059555064724
‘setupmefree’ 0.9966181271429526
‘com’ 0.3984979685547945
‘http’ 0.1447738056917774
‘a’ 0.06584106490791478
Table 8: The positive and negative weights using LIME for second text
Positive Numbers Negative Numbers
Word Value Word Value
‘megapixel’ 0.45124215111245936 ‘a’ -0.09469945665856638
‘get’ 0.4369880334209157
‘free’ 0.4206519368628723
‘digital’ 0.3721259726068145
‘camera’ 0.2388125335048898
‘canon’ 0.19498844421129996
‘6’ 0.04953986530836751
‘3’ 0.024635744735869915
Along with LIME, we also examined the model interpretability with similar text by using Transformer Interpret. During this interpretation, this Transformer Interpret uses the tokenizer to split the large words into sub-words with the ‘##’ prefix. This happens when an unusual or non-vocabulary word is encountered by the model, which breaks it down into smaller sub-words to enable efficient processing. Fig. 9 shows Transformer Interpretation for a phishing email and Table  9 and Table  10 show the word importance, with positive and negative numbers associated with each word for two phishing emails.
(a) Explanation for first text
(b) Explanation for second text
Figure 9: Transformer Interpret Explanation for two Phishing Email Text
Table 9: The positive and negative numbers using Transformer Interpret for first Phishing Email Text
Positive Numbers Negative Numbers
Word Value Word Value
‘low’ 0.7698059423920272 ‘.’ -0.09537831853568463
‘price’ 0.4465021887746993 ‘setup’ -0.037898186884516454
‘software’ 0.10363795047092349 ‘?’ -0.04837875144898304
‘http’ 0.05008396373925551 ‘a’ -0.212114259929307
‘:’ 0.06724251662885174 ’=’ -0.13427331689723981
‘/’ 0.018869178109329452 ‘310’ -0.06162690120821821
‘neon’ 0.07499325538696087 ‘##7’ -0.013360917838605864
‘##ate’ 0.09833698869154901 ‘.’ -0.1269640647507237
‘me’ 0.09196861346171359
‘free’ 0.16949987141093556
‘com’ 0.20340796716052334
‘/’ 0.006729002810807768
Table 10: The positive and negative numbers using Transformer Interpret for second Phishing Email Text
Positive Numbers Negative Numbers
Word Value Word Value
‘get’ 0.49733595483460524 ‘a’ -0.2688505206057748
‘free’ 0.4645088664478948 ‘6’ -0.09532782298942542
‘mega’ 0.18191195439136695 ‘.’ -0.08192496249808812
‘3’ 0.014327378740528552 ‘##pi’ -0.11609630028309895
‘##x’ 0.5855798135703232 ‘!’ -0.18838964048006548
‘digital’ 0.021290989158880363
‘canon’ 0.13188078942699152
‘camera’ 0.05155301993306188
‘>’ 0.03909299350598931
In Fig. 9 the highlighted text with green color shows the word importance to classify the analyzed text as Phishing. On the other hand, the red portions showed those words or sub-words that worked against the classified class. As we worked with a binary classification task here “True Label" = 1 means "Phishing Email" and when “True Label" = 0 means “Safe Email”.
@@
According to the analysis, both explainable approaches explain how the suggested model predicts "Phishing Email" for each of the two tested input texts, but they differ in how they define explanation. LIME mostly uses small textual modifications to evaluate the impact of the model’s prediction on the entire text, while Transformer Interpret employs an attention mechanism to examine each segment of the input sequence. Furthermore, in the LIME and Transformer Interpret explanation processes, the resulting quantity of tokens, words, or sub-words differs for the two phishing email contents. The positive words are mainly responsible for the predicting class. More specifically, when those words are present in a text that time model predicts a class. A comprehensive difference between the two explainable AI techniques is shown in Fig. 10 and Fig. 11. The number of scoring percentages for those email texts is not the same for LIME and Transformer Interpret. This is only because the working procedures of those techniques are not the same.
(a) LIME Scoring
(b) Transformer Interpret Scoring
Figure 10: LIME and Transformer Interpret Comparison on first text
(a) LIME Scoring
(b) Transformer Interpret Scoring
Figure 11: LIME and Transformer Interpret Comparison on second text
@@ 6 Discussion
The results of both an imbalanced and balanced dataset depict an enhancement in the performance of the proposed transformer-based fine-tuned DistilBERT model. The fine-tuned version of the DistilBERT model showed 98.90% training and 97.50% testing accuracy on an imbalanced dataset whereas this model showed 99.07% training and 98.48% testing accuracy on the balanced dataset. Although these results show that the model’s performance on a balanced dataset is quite better than an imbalanced dataset, the difference is not so high, approximately 1%. This demonstrated that the fine-tuned version performs well in an imbalanced dataset. Additionally, the performance shows that the optimized model performed better than other previous studies.
Besides, this research work also answered the “How was the model performed?" and “Why was it performed?" using XAI, motivated from recently published book by Sarker (Sarker, 2024). Two well-known Explainable and interpretable techniques are utilized for the model understanding. It is more difficult to understand how DistilBERT makes its predictions because, as we know, it is a complex model. These AI techniques aided in gaining an understanding of the model’s decision-making process. Both LIME and Transformer Interpret highlighted the areas of the text that are more responsible for the classification of the phishing of safe email. This highlights areas worked with words or sub-words, making the decision more understandable for humans, and users can also comprehend it easily. Mainly, those techniques calculated each word or sub-words numeric value or weight where some words indicate positive value and some negative. Words with calculated higher positive values are more important for classifying the predicted class, while those with more negative values do not support the predicted class. This experimental analysis showed the transparency of the model in this phishing email classification task.
@@ 7 Conclusion
This study has investigated the detection of phishing emails by utilizing a Large Language Model. This also has shown how effective it is to use state-of-the-art natural language processing methods to improve the precision and effectiveness of phishing email identification. The overall experiment was done with both a balanced and imbalanced dataset. We have obtained promising results in differentiating between phishing and valid emails by utilizing DistilBERT’s contextual language understanding technique. Mainly, to achieve this good performance the model is optimized by changing some important parameters such as batch size, optimization function, learning rate, etc, and also fine-tuning has been done. This fine-tuned version of DistilBERT obtained 98.48% testing accuracy in this phishing email classification task. Furthermore, LIME and Transformer Interpret techniques have been employed to see the explanation of the model and how it makes decisions. The decision-making process is now more visible and understandable for users because of the usage of explainable AI, which promotes comprehension and trust. The overall model explainability describes the robustness of the proposed model.
The field of Large Language Models has experienced quick and dynamic development due to the substantial investment made by both businesses and consumers. New open-source models are being produced almost every day, demonstrating the ongoing development and innovation in this field. In the future, we plan to expand the scope of our research by experimenting with a wider variety of models and datasets. Furthermore, we intend to investigate the use of different variants of BERT-based models for the identification of different language phishing emails. Moreover, deep analysis of the model will be conducted to ensure its robustness by enhancing the explainable AI techniques.
@@
The dataset that is used in this study is available at Kaggle (Online Link):  https://www.kaggle.com/datasets/subhajournal/phishingemails.

### v2

@@ 5 Explainability Analysis
LIME and Transformers Interpret both demonstrated how the fine-tuned RoBERTa model arrives at its prediction for phishing email detection by approximating the model’s behavior on every word of the email text. These explainability methods provided accurate results by answering “Why” the model classifies a selected text as a legitimate or phishing email. Based on the explainability, these methods assign positive and negative importance scores to each word, indicating their contribution to the predicted class. These scores highlight which word was more specifically reasonable for the predicted class. By using those scores, the explainability models visualize the whole text by highlighting the words that scored positive or negative on the prediction.
Fig. 9 shows the LIME-based explanation for two phishing email samples, which the model also classified as phishing. The contributions can be seen in context by highlighting keywords in the text instances with the appropriate color. To better understand the explanation produced by the LIME explainer, presented bar plots of those phishing email samples by indicating the top 15 words ranked by their contribution to the classification. The bar length with weight indicates the strength of each word. Furthermore, to make the explanation relevant to the most likely result, we concentrated on the top-predicted label. This is especially helpful when the main goal is to comprehend the choice for the most likely class.
(a) Explanation for the phishing sample 1
(b) Explanation for the phishing sample 2
Figure 9: LIME Explanation for two Phishing Email Samples
In addition to LIME, we used Transformers Interpret to analyze model interpretability through comparable phishing examples. This tool examines how each token influences the model’s final prediction to generate token-level attribution scores. Since RoBERTa uses a BPE tokenizer, complex or rare words are broken into smaller word units, making it easier for the model to handle out-of-vocabulary terms. Figure 10 illustrates a visual interpretation for phishing email samples, where tokens highlighted in green contributed positively to the classification as phishing email, and tokens in red contributed negatively.
(a) Explanation for sample 1
(b) Explanation for sample 2
Figure 10: Transformers Interpret Explanation for two Phishing Email Samples
@@
According to the analysis, both explainable approaches explain how the suggested model predicts "Phishing Email" for each of the tested input samples. However, those techniques differ fundamentally in how the explanations are generated. LIME primarily uses small textual perturbations to evaluate the impact of individual words on the model’s prediction across the entire text. At the same time, Transformers Interpret employs a gradient-based explanation technique to analyze the contribution of each token or sub-word to the model’s final prediction. Furthermore, in the LIME and Transformers Interpret explanation processes, the resulting quantity of tokens, words, or sub-words differs for the two phishing email samples. In both techniques, the positively contributing words are mainly responsible for the predicted class. More specifically, the presence of those tokens significantly increases the likelihood of the model predicting the sample as phishing.
In our study, we found that both methods can have opposite polarities for the same words. For example, one explainable method finds a word to be positive and the other negative, which can lead to confusion in the mind of end-users. To address this issue, we developed our hybrid approach, LITA, for our analysis. By integrating the complementary strengths of the two methods, we achieve a more balanced explanation. The LITA also generated explanations for two phishing samples shown in Fig. 11; and Table 11 and Table 12 present the top 15 words ranked by their LIME scores for each sample, along with the corresponding scores from Transformers Interpret and the hybrid LITA approach. These tables present quantitative evidence that highlights interpretability differences between both individual methods and illuminates how LITA reconciles aligned and conflicting attributions to improve clarity and trust.
(a) Explanation for the phishing sample 1
(b) Explanation for the phishing sample 2
Figure 11: LITA Explanation for two Phishing Email Samples
Table 11: Word-Level Explainability Analysis for Phishing Sample 1 Using LIME, Transformers Interpret, and Hybrid Scoring
Word LIME Score LIME (Norm) Score Transformer Score Transformer (Norm) Score Hybrid Score
freida 0.35498 1.00000 0.10828 0.45581 0.72791
verify 0.32802 0.91895 -0.01591 -0.32235 0.29830
senior -0.31029 -1.00000 0.00691 -0.17936 -0.58968
confirm 0.29308 0.81391 -0.06232 -0.61315 0.10038
loan 0.27749 0.76704 0.02076 -0.09258 0.33723
here 0.25293 0.69321 0.00004 -0.22241 0.23540
webber 0.25283 0.69291 0.16803 0.83020 0.76155
applicant -0.24042 -0.78995 -0.12406 -1.00000 -0.89497
application -0.21336 -0.70860 -0.08878 -0.77894 -0.74377
a -0.19875 -0.66468 0.02981 -0.03587 -0.35027
financial 0.19619 0.52263 0.08843 0.33143 0.42703
000 0.18949 0.50249 0.10274 0.42110 0.46179
association 0.18149 0.47844 0.19513 1.00000 0.73922
your 0.17999 0.47393 -0.00322 -0.24283 0.11555
manager -0.17817 -0.60281 -0.01086 -0.29070 -0.44676
Table 12: Word-Level Explainability Analysis for Phishing Sample 2 Using LIME, Transformers Interpret, and Hybrid Scoring
Word LIME Score LIME (Norm) Score Transformer Score Transformer (Norm) Score Hybrid Score
free 0.18759 1.00000 0.13196 0.39400 0.69700
perfect 0.14421 0.64174 0.15117 0.46549 0.55361
removed 0.10480 0.31627 0.18932 0.60745 0.46186
want 0.09803 0.26035 0.20522 0.66662 0.46349
remove 0.08880 0.18413 0.13166 0.39289 0.28851
special 0.08735 0.17215 0.13495 0.40513 0.28864
no 0.06542 -0.00896 -0.21209 -0.88628 -0.44762
type 0.05757 -0.07379 0.01917 -0.02571 -0.04975
somebody 0.05757 -0.07379 0.00061 -0.09478 -0.08429
the -0.05458 -1.00000 0.03316 0.02635 -0.48683
gift 0.05245 -0.11608 0.29481 1.00000 0.44196
be 0.04975 -0.13837 0.00989 -0.06025 -0.09931
further 0.04455 -0.18132 -0.24265 -1.00000 -0.59066
catalog -0.04401 -0.91271 -0.21059 -0.88070 -0.89670
please -0.03814 -0.86423 0.02428 -0.00670 -0.43546
@@ 6 Discussion
The results of both an imbalanced and a balanced dataset depict an enhancement in the performance of the proposed transformer-based fine-tuned RoBERTa model. The fine-tuned version of the RoBERTa model showed 98.57% training and 97.96% testing accuracy on an imbalanced dataset, whereas this model showed 98.94% training and 98.45% testing accuracy on the balanced dataset. Although these results show that the model’s performance on a balanced dataset is quite better than on an imbalanced dataset, the difference is not so high, approximately 0.49%. This demonstrated that the fine-tuned version performs well in an imbalanced dataset. Additionally, the performance shows that the optimized model performed better than other previous studies.
Additionally, this research work also addressed the questions “How was the LLM model performed?" and “Why was it performed?" using XAI, as motivated by the recently published paper by Sarker (Sarker, 2024). The author highlighted the potentiality of the LLMs by addressing the technical, ethical, and societal challenges, focusing on fairness, transparency, transparency, explainability, and trust. By summarizing and understanding this potential research scope with direction, two well-known explainable and interpretable techniques are utilized in our research to understand the model’s decision. It is more difficult to understand how RoBERTa makes its predictions because, as we know, it is a complex model. These AI techniques aided in gaining an understanding of the model’s decision-making process. Both LIME and Transformers Interpret highlighted the areas of the text that are more responsible for the classification of phishing or legitimate emails. This highlights areas worked with words or sub-words, making the decision more understandable for humans, and users can also comprehend it easily. Mainly, those techniques calculated each word or sub-words numeric value or weight, where some words indicate positive value and some negative. Words with calculated higher positive values are more important for classifying the predicted class, while those with more negative values do not support the predicted class. This experimental analysis showed the transparency of the model in this phishing email classification task. Although those methods show almost similar explanations, but in some word cases their explanation not agreed due to their different functionality. But this can create serious confusion which one solved in this work, with LITA.
@@ 7 Conclusion
This study explores the detection of phishing emails using a transformer-based language model. It also demonstrates the effectiveness of employing state-of-the-art natural language processing methods to improve the accuracy and efficiency of identifying phishing emails. The overall experiment was conducted with both balanced and imbalanced datasets. We achieved promising results in distinguishing between phishing and legitimate emails by utilizing RoBERTa’s contextual language understanding capabilities. To accomplish this, the model was optimized by adjusting key parameters such as batch size, optimization function, learning rate, and through fine-tuning. This fine-tuned version of RoBERTa obtained a 98.45% testing accuracy in the phishing email classification task. Additionally, to improve interpretability and foster greater trust in the model’s explanations, we introduced LITA, a hybrid explanation approach that combines LIME and Transformers Interpret. This makes the decision-making process more transparent and understandable for users, promoting comprehension and trust. Overall, the model’s explainability reflects the robustness of the proposed system.
The field of LLMs has experienced quick and dynamic development due to the substantial investment made by both businesses and consumers. New open-source models are being produced almost every day, demonstrating the ongoing development and innovation in this field. In future work, we plan to explore various LLMs, including LLaMA, Gemini, Falcon, Mistral, and others, to assess their effectiveness in phishing email detection. We will extend our research to multilingual datasets to improve our model’s performance across different languages and cultural contexts. To further enhance the robustness and accuracy of our algorithms, we also plan to create a new dataset comprising phishing emails. Additionally, we aim to integrate our models into real-world cybersecurity systems, partnering with industry players to evaluate their real-time effectiveness and refine them based on practical feedback. Furthermore, we will conduct more in-depth analysis of the models’ decision-making processes to ensure robustness by advancing explainable AI techniques.
@@
The research dataset which one is used in this study is available at Kaggle:  https://www.kaggle.com/datasets/subhajournal/phishingemails.